OpenStudy (anonymous):

A bullet is shot at a target that is 30.0 m away. If it has a muzzle velocity of 300 m/s, the time it will take to hit the target is

OpenStudy (anonymous):

neglecting losses the distance to be travelled ie 30m is 1 tenth of the distance the bullit could travel in 1 second.

OpenStudy (anonymous):

so 1 second or 0.1 second?

OpenStudy (anonymous):

Good job its 0.1 seconds. 1/10=0.1

OpenStudy (anonymous):

can i ask how experienced you are in this material

OpenStudy (anonymous):

I have a degree in mechanical engineering

OpenStudy (anonymous):

awesome. if i type out the full problem could you help me with this assignment?

OpenStudy (anonymous):

Perhaps if it is not long I must go to bed soon its very late

OpenStudy (anonymous):

okay just tell me when youre going to go to bed

OpenStudy (anonymous):

Jose can jump vertically 1 meter from his skateboard when it is at rest. When the skateboard is moving horizontally with a speed of 2 m/s, Jose can jump a) the same b) less than 1 m c) more than 1 m

OpenStudy (anonymous):

Horizontal and vertical motion can be considered independently so a)the same Although from personal experience it is more scary to jump on a skate board when moving.

OpenStudy (anonymous):

mvery scary haha A cannonball is fired at some angle in the air. In the first second it moves 10.0 m horizontally. Assuming it doesnt hit the ground and if air resistance is ignored, how far will it move horizontally in the next second? a) more than 10 m b) 10 m c) less than 10 m d) not enough info to say

OpenStudy (anonymous):

Similar to before since there is no air resistance there are no losses so the canon ball will continue in exactly the same horizontal distance in the subsequent seconds as it did in the first. until it hits the floor

OpenStudy (anonymous):

one more then i must go

OpenStudy (anonymous):

okay Cletus is trying to throw a rope to a boat that is stranded off shore. if he is standing on a ship that is 3.5 meters above the boat, how fast must he throw the rope horizontally to make sure it will reach the boat that is 10 m away?

OpenStudy (anonymous):

Gravity =10 m/s. since acceleration = distance/time squared. consider the vertical distance the rope can fall which is 3.5m. The time taken to fall that far is given t=sqroot(3.5/10). So it takes 5.9 seconds to fall the 3.5m. Now considering the horizontal distance of 10m which has to be covered in 5.9sec leads to..... velocity in x direction = distance in x direction / time to cover distance ie 10/5.9=16 m/s

OpenStudy (anonymous):

Hope you find that helpful :)

OpenStudy (anonymous):

eddit gravity = 10 m/s/s

OpenStudy (anonymous):

thanks you so much!

OpenStudy (anonymous):

Good night